100μL of 1M sodium chloride (NaCl) is mixed with100μL of water into a single tube.

What is the final concentration of sodium chloride?
Answer #1

Solution:-

Initial Concentration of NaCl solution = 1M,

Initial volume of NaCl solution is = 100 × 10-6 L

Mass of NaCl dissolved (w) = ?

Molarity (M) = no. of moles of NaCl / Volume in Litres

no. of moles of NaCl = Mass of NaCl ( in grams) dissolved / Molar Mass of NaCl

= w ( in gm) / 58.5

Thus, using above information,

1 = w / 58.5 × 100× 10-6

w = 1 × 58.5 × 100× 10-6

= 0.00585 grams

Now, 100 × 10-6 L more water is added in NaCl solution

Final volume of NaCl solution = 100 × 10-6 + 100 × 10-6

= 200 × 10-6 L

​​​​​​w = 0.00585 grams

no. of moles of NaCl = 0.00585/ 58.5

= 0.0001

Molarity = no.of moles of NaCl / Final Volume of NaCl

= 0.0001/ 200 × 10-6

= 0.5 M

Final concentration of NaCl solution is = 0.5 M